write the equation of the line in slope-intercept form
passing through the points (-1,5)(-3,-2)

Answer by checkley77(12844)

(-1,5)(-3,-2)
SLOPE=(-2-5)/(-3+1)
SLOPE=-7/-2
SLOPE=7/2
Y=mX+b
5=7/2*-1+b
5=-7/2+b
b=5+7/2
b=(10+7)/2
b=17/2 THE Y INTERCEPT.
Y=7X/2+17/2